# A quick probability-oriented introduction to operator splitting methods

### v1

@@ 1. The scope of the paper and first examples
where $A$ represents diffusion, the field $B$ is (possibly superficial) velocity and $R$ is sources and sinks. $R$ may be of chemical origin, while the velocity field reflects physical properties of a reservoir etc. One may prefer drastically different numerical methods for integrating subsystems obtaining by splitting the RHS into a sum of differential operators of at most second order and treating those subsystems separately. For instance, physically justified first order equations have a developed theory that often explicitly relies on conservation laws (e.g. [188]) while the diffusion part is typically treated via finite elements (e.g. [351]).
@@
can be solved explicitly. This is a toy model example where a stiff ($A_{1}$) and a non-stiff part ($A_{2}$) of the original equation can be isolated. Henceforward stiffness (e.g. [168, 226, 188] in the deterministic setting and [253] in the stochastic setting) in the context of operator splitting methods is understood in a vague general sense: a system is stiff if a successful application of certain numerical methods (usually explicit ones and/or those used in the field as standard integrators) requires an impractically small time step and is thus infeasible or costly/hard to implement. The idea of the current example combines well with the previous example since competing mechanisms of different origin (chemical vs. mechanical) may act on different scales, which is a typical source of stiffness as it directly increases the stiffness ratio.
@@
(e.g. [40]) As a development of Example 2, consider a stiff ODE in ${\mathbb{R}}^{d}$ with constant coefficients
@@
Combining operators $(T^{(1)}_{t})$ and $(T^{(2)}_{t})$ on the interval $[0;n^{-1}]$ gives for any $x=(x_{1},x_{2})$
@@
As a result, operator splitting procedures are often developed for equations and problems in a general formulation (for instance, for the Navier-Stokes equation, the Zakai equation of the filtration theory, the Allen-Kahn equation as an example of the advection-diffusion-reaction system and composite optimization problems as an example in a non-PDE related setting).
@@
  • to provide a wider picture and to at least mention directions the whole theory goes beyond the Trotter-Kato formula in the context of semigroups;
  • to give examples of results developed through the use of operator splitting methods in the field of probability and to show that randomness provide new insights and may lead to – and require, in fact, – the development of new techniques.
@@
At the same, this also leads to some asymmetries in the text: the discussion of the BCH formula is rather lengthy only because such material is not expected to covered in textbooks on probability while the reader is assumed to be familiar with SPDEs and variational methods for them so the part devoted to SPDEs does not explain the setting.
It should be emphasized that there is more than enough information around for a person interested in the topic, including high quality introductory level texts and textbooks. Not even trying to compete with classical texts, [^4] There are also highly efficient surveys (e.g. [88, 264, 42]). we merely concentrate on gathering (at least in the form of bibliographic references) important basic facts, hints and probabilistic results as expanded lecture notes (for the aforementioned course) and giving a total newcomer basic understanding and a map to navigate.
@@
  • [280, 151, 121, 94, 200, 77, 122] for the Trotter-Kato formula, the perturbation theory for semigroups and the formulation in the terms of abstract evolution equations,
  • [148, 126] [^5] Chapters 1–2 for a general exposition and a survey. for a survey, the general theory, history, connections to optimization problems, PDEs and variational inequalities, and examples,
  • [188] for a treatment of advection-diffusion-reaction equations, examples and a discussion of numerics,
  • [289, 308, 205, 240, 171, 290, 346] for the physical exposition and applications to Feynman path integrals and the Feynman-Kac formula,
  • [185] for operator splitting methods for rough solutions,
  • [148, 150, 149, 340, 294, 28, 246, 247, 248] for surveys, innumerable references, applications, and using operator splitting numerical methods for solving PDEs and variational nonlinear inequalities, (convex) optimization problems (including sparse and large scale problems), fixed point algorithms, problems for monotone operators etc.,
  • [169, 40, 252, 297] for applications to geometric integration of ODEs.
@@ 2. Matrices and the Lie product formula
Theorem 1 (e.g. [289, 171]).
@@
The original source for Theorem 1 seems to be untrackable [80].
@@
which finishes the proof [^7] We get the first order for the whole method and not the second order since (2.3) is a local error which gets propagated exactly $n$ times. This is a universal rule (cf. [253])..
@@
Theorem 1 also finds applications in the theory of matrix Lie groups and the corresponding homeomorphisms [170], matrix trace inequalities [282] etc.
The next theorem was originally proposed for one two-dimensional difference scheme and can also be proved with direct calculations.
Theorem 2 (e.g. [312]).
@@
The local error of the Strang splitting is [188]
@@ 3. Semigroups and the Chernoff product formula for abstract Cauchy problems
To venture forth, we need to recall standard material on semigroups and the abstract Cauchy problem (ACP), the Trotter-Kato theorems and the Chernoff product formula [280, 151, 121, 94, 200, 77, 122]. We omit some details and technicalities. All integrals are Bochner integrals.
@@
where $C_{0}({\mathbb{R}}^{m})$ is the space of continuous function that vanish at infinity [122, 50] [^8] The very definition of a Feller process may vary between authors [50].. In particular, Feller’s one-dimensional diffusion on an interval $I$ has generator
@@
while the precise description of $D(A)$ incorporates boundary conditions (and thus the behavior of the process at the boundary) and $c$ corresponds to killing. Note that it is not possible in general to give a full description of $D(A)$ for a Feller process and it is often sufficient to consider a core $L$ instead, that is, a set $L\subset D(A)$ such that the closure of $A|_{L}$ equals $A$ [122].
@@
A Wiener process analogously defines a $C_{0}-$semigroup on $L_{2}({\mathbb{R}}^{m})$ with generator $\frac{1}{2}\Delta$ and $D(A)=W^{2,2}({\mathbb{R}}^{m})$ [44]. However, the possibility to extend a Feller semigroup onto some $L_{p}$ is linked to the properties of the adjoint operator and thus to properties of Fokker-Plank-Kolmogorov equations (for discussions and conditions see e.g. [313, 72, 50] and [121, Section VI.4]).
@@
is called an evolution problem. A classical solution is defined similarly to the autonomous case but no simple analog of Theorem 3 exists (see [280, 151] for the theory and [259, 301] and [25, Remark 1.12] for a survey and references). Moreover, (3) requires the compatibility of $\{D(A(t))\mid t\in[0;T]\}$ if the domains in question are time-dependent, a standard condition being $\overline{\cap_{t\in[0;T]}D(A(t))}=X.$ The equation (3) is then solved for $x\in\cap_{t\in[0;T]}D(A(t)).$
@@
In general, non-autonomous inhomogeneous ACPs, unavoidable in applications, may behave in an unexpected way and one should be careful to draw conclusions about the properties of the corresponding propagators [259].
@@
The following theorem is a version of the celebrated Trotter-Kato theorem999occasionally called the Trotter-Neveu-Kato theorem (e.g [344, 68, 43]) whose general form provides a basis for numerous approximation schemes in the theory of semigroups, including Yosida approximations [^10] For the exposition of SPDEs in the terms of Yosida approximations see [161].. One can consult [154, 153, 200] for a survey of the general approximation theory of $C_{0}-$ semigroups, some results about the rate of the convergence (in the operator norm, too) and further references. A number of references concerning the Chernoff and Trotter-Kato product formulas will be given later.
@@
[283, 285, 286, 284] use probabilistic methods to develop a unified exposition of many approximation formulas (see also [33]).
@@
it holds that (1) normal-$\Rightarrow$ (2) and (2) normal-$\Leftrightarrow$ (3).
@@
A probabilistic formulation of Theorem 4 for Feller processes can be found in [206]. Trotter-Kato theorems are behind some results about approximations of Feller processes with Markov chains, in particular [206].
@@
Counterexamples for the Trotter-Kato theorem, in particular of probabilistic origin, can be found in [43]; variational and nonlinear versions of the Trotter-Kato theorem, in [200]; generalizations for norm operator topology, in [344, 68].
The next theorem is called the Chernoff product formula and is another fundamental tool in the approximation theory of $C_{0}-$semigroups.
@@
The formula (3.7) provides an approximation of the semigroup $(T_{t})_{t\geq 0}$ in the terms of the family $V.$ The exposition of the method of obtaining such approximations is given in a survey [63]; further references and applications to stochastic processes including those on manifolds and in domains with Dirichlet and Robin boundary conditions can be found in [63, 61, 60, 251, 64, 271, 303, 269, 311, 251, 273, 309, 62, 65, 64, 61, 310]. Other references given later in the text for the Trotter-Kato theorem in the context of the approximation theory are relevant, too.
The following additional references illustrate a possibility to strengthen the conclusion of the Chernoff product formula. [344, 343, 345, 68, 138, 139, 288] establish convergence in the operator norm an/or obtain estimates of the rate of the convergence for self-adjoint operators and quasi-sectorial contraction semigroups etc.; [279, 343, 344, 68] obtain estimates by using probabilistic arguments. It is worth noting that the convergence can be arbitrary slow and may not hold in a stronger topology.
@@
([63]; cf. [229]) Assume that real-valued $a,\sigma$ functions are Lipschitz continuous and bounded, $\inf_{u\in{\mathbb{R}}}|\sigma(u)|>0$ and consider a SDE
@@
for sufficiently smooth compactly supported $f.$ See also [311, 251, 310] for other applications of the Chernoff product formula to pathwise approximations of random processes. Note that this result does not recover the first order of the weak Euler-Maruyama scheme.
@@
([64]) Let $g_{\mu}$ be the one-dimensional Gaussian density with 00 mean and variance $\mu.$ (3.10) can be rewritten as
@@
Such a representation of the original semigroup as the limit of iterated integrals w.r.t. some finite-dimensional projections of a (pseudo)measure on a phase space (the Wiener measure on $C([0;T])$ here) is called the Feynman formula [309] [^11] The definition of the Feynman formula may slightly vary (c.f. [65]). and presents an alternative approach to the Feynman-Kac formula. See [63, 60, 311, 251, 273, 309, 62, 65, 64, 61, 310] for other results on Feynman formulas.
@@
and passing to the limit in (3.10) and (12) gives a particular case of the Girsanov theorem (e.g. [234]) for $x$ and $y:$
@@
The behavior of a solution of an PDE and its differentiability at $t=0$ (and thus the behavior of the corresponding semigroup) may a be subtle moment in applications, particularly when a probabilistic interpretation is used. A typical example is the function $u(x,t)=\operatorname{E}_{x}f(\xi(\min(t,\tau))),x\in D,t\geq 0,$ where $\tau$ is the moment a Feller process $\xi$ hits the boundary of a domain $D.$ If $f=\!\!\,{\textbf{1}}_{\partial D}$ then $u=\operatorname{P}(\tau<t)$ is discontinuous at $t=0$ though $u\in C^{2,1}(D^{int}\times(0;\infty))$ usually. As a result, the weak variational formulation in the terms of Gelfand triples [278] or the weak distributional formulation [230, 325] can be suggested as an alternative (see also [129, 335]).
@@
One particular example of the versatility of the Chernoff product formula is the possibility to use it to prove the central limit theorem [151] .

### v2

@@ 1. The scope of the paper and first examples
where $A$ represents diffusion, the field $B$ is (possibly superficial) velocity and $R$ is sources and sinks. $R$ may be of chemical origin, while the velocity field reflects physical properties of a reservoir etc. One may prefer drastically different numerical methods for integrating subsystems obtaining by splitting the RHS into a sum of differential operators of at most second order and treating those subsystems separately. For instance, physically justified first order equations have a developed theory that often explicitly relies on conservation laws (e.g. [HundsVer07Numerical]) while the diffusion part is typically treated via finite elements (e.g. [ZienTayZhu13finite]).
@@
can be solved explicitly. This is a toy model example where a stiff ($A_{1}$) and a non-stiff part ($A_{2}$) of the original equation can be isolated. Here stiffness (e.g. [HaiWan96solving, Lam91numerical, HundsVer07Numerical] in the deterministic setting and [Mil95numerical] in the stochastic setting) in the context of operator splitting methods is understood in a vague general sense: a system is stiff if a successful application of certain numerical methods (usually explicit ones and/or those used in the field as standard integrators) requires an impractically small time step and is thus infeasible or costly/hard to implement. The idea of the current example combines well with the previous example since competing mechanisms of different origin (chemical vs. mechanical) may act on different scales, which is a typical source of stiffness as it directly increases the stiffness ratio.
@@
(e.g. [BlaCa16concise]) As a development of Example 2, consider a stiff ODE in ${\mathbb{R}}^{d}$ with constant coefficients
@@
Combining operators $T^{(1)}_{t}$ and $T^{(2)}_{t}$ on the interval $[0;n^{-1}]$ gives for any $x=(x_{1},x_{2})$
@@
As a result, operator splitting procedures are often developed for equations and problems in a general formulation (for instance, for the Navier-Stokes equation, the Zakai equation of the filtration theory, advection-diffusion-reaction equations, and composite optimization problems as an example in a non-PDE setting).
@@
  • to provide a wider picture and to at least mention directions the whole theory goes beyond the Trotter-Kato formula;
  • to give examples of results developed through the use of operator splitting methods in the field of probability and to show that randomness leads to new insights and new techniques;
@@
At the same, this also leads to some asymmetries in the text: the discussion of the Baker-Campbell-Hausdorff formula is rather lengthy only because such material is not expected to be covered in textbooks on probability while the reader is assumed to be familiar with SPDEs and variational methods for them so the part devoted to SPDEs does not explain the setting.
It should be emphasized that there is more than enough information around for a person interested in the topic, including high quality introductory level texts and textbooks. Not even trying to compete with classical texts, [^4] There are also highly efficient surveys (e.g. [CsoSi21numerical, NeidSteZa18operator, BlaCaMu08splitting]). we merely concentrate on gathering (at least in the form of bibliographic references) important basic facts, hints and probabilistic results as expanded lecture notes (for the aforementioned course) and giving a total newcomer basic understanding and a map to navigate.
@@
  • [Pazy12Semigroups, Goldstein85semigroups, EnNa00One-parameter, Da0One-parameter, ItoKa02evolution, Cher74product, EthKur09Markov] for the Trotter-Kato formula, the perturbation theory for semigroups and the formulation in the terms of abstract evolution equations,
  • [GlowOshYin17Splitting, FaHa09splitting] [^5] Chapters 1–2 for a general exposition and a survey. for a survey, the general theory, history, connections to optimization problems, PDEs and variational inequalities, and examples,
  • [HundsVer07Numerical] for a treatment of advection-diffusion-reaction equations, examples and a discussion of numerics,
  • [ReedSai72methods, Si05functional, JohnLa00Feynman, LoHiBetz20FeynmanKac, Hall13quantum, ReedSi75methods_2, Zeid95applied] for the physical exposition and applications to Feynman path integrals and the Feynman-Kac formula,
  • [HolKarlKenLie10Splitting] for operator splitting methods for rough solutions,
  • [GlowOshYin17Splitting, GloLeTa89Augmented, Glo03Finite, Ya71Method, RyuYin23large_scale, BauCom17convex, Mar82methods, Mar88splitting, Mar86splitting] for surveys, innumerable references, applications, and using operator splitting numerical methods for solving PDEs and variational nonlinear inequalities, (convex) optimization problems (including sparse and large scale problems), fixed point algorithms, problems for monotone operators etc.,
  • [HaiLu10geometric, BlaCa16concise, McLachQuis02splitting, SanzCal94numerical] for applications to geometric integration of ODEs.
@@ 2. Matrices and the Lie product formula
Theorem 1 (e.g. [ReedSai72methods, Hall13quantum]).
@@
The original source for Theorem 1 seems to be untrackable [CoFriedKaKe82eigenvalue].
@@
which finishes the proof [^7] We get the first order for the whole method since (2.3) is a local error which gets propagated exactly $n$ times. This is a universal rule (cf. [Mil95numerical])..
@@
Theorem 1 also finds applications in the theory of matrix Lie groups and the corresponding homeomorphisms [Hall15Lie], matrix trace inequalities [Petz94survey] etc.
The next theorem was originally proposed for one two-dimensional difference scheme and can also be proved by direct calculations.
Theorem 2 (e.g. [Strang68construction]).
@@
The local error of the Strang splitting is [HundsVer07Numerical]
@@ 3. Semigroups and the Chernoff product formula for abstract Cauchy problems
To venture forth, we need to recall standard material on semigroups and the abstract Cauchy problem (ACP), the Trotter-Kato theorems and the Chernoff product formula [Pazy12Semigroups, Goldstein85semigroups, EnNa00One-parameter, Da0One-parameter, ItoKa02evolution, Cher74product, EthKur09Markov]. We omit some details and technicalities. All integrals are Bochner integrals.
@@
where $C_{0}({\mathbb{R}}^{m})$ is the space of continuous function that vanish at infinity [EthKur09Markov, BottSchiWang13Levy] [^8] The very definition of a Feller process may vary between authors [BottSchiWang13Levy].. In particular, Feller’s one-dimensional diffusion on an interval $I$ has generator
@@
while the precise description of $D(A)$ incorporates boundary conditions (and thus the behavior of the process at the boundary) and $c$ corresponds to killing. Note that it is not possible in general to give a full description of $D(A)$ for a Feller process and it is often sufficient to consider a core $L$ instead, that is, a set $L\subset D(A)$ such that the closure of $A|_{L}$ equals $A$ [EthKur09Markov].
@@
A Wiener process analogously defines a $C_{0}-$semigroup on $L_{2}({\mathbb{R}}^{m})$ with generator $\frac{1}{2}\Delta$ and $D(A)=W^{2,2}({\mathbb{R}}^{m})$ [BoKryRo15Fokker]. However, the possibility to extend a Feller semigroup onto some $L_{p}$ is linked to the properties of the adjoint operator and thus to properties of Fokker-Plank-Kolmogorov equations (for discussions and conditions see e.g. [Stroock2008Partial, Cas85generators, BottSchiWang13Levy] and [EnNa00One-parameter, Section VI.4]).
@@
is called an evolution problem. A classical solution is defined similarly to the autonomous case but no simple analog of Theorem 3 exists (see [Pazy12Semigroups, Goldstein85semigroups] for the theory and [NaNi02well-posedness, Schnau02wellposedness] and [BatCsoFarNi11operator, Remark 1.12] for a survey and references). Moreover, (3) requires the compatibility of $\{D(A(t))\mid t\in[0;T]\}$ if the domains in question are time-dependent, a standard condition being $\overline{\cap_{t\in[0;T]}D(A(t))}=X.$ The equation (3) is then solved for $x\in\cap_{t\in[0;T]}D(A(t)).$
@@
In general, non-autonomous inhomogeneous ACPs, unavoidable in applications, may behave in an unexpected way and one should be careful to draw conclusions about the properties of the corresponding propagators [NaNi02well-posedness].
@@
The following theorem is a version of the celebrated Trotter-Kato theorem999occasionally called the Trotter-Neveu-Kato theorem (e.g [Zag22notes, CaZag01operator, Bob07limitations]) whose general form provides a basis for numerous approximation schemes in the theory of semigroups, including Yosida approximations [^10] For the exposition of SPDEs in the terms of Yosida approximations see [Go16Yosida].. One can consult [GoTo14convergence, GoKoTo19general, ItoKa02evolution] for a survey of the general approximation theory of $C_{0}-$ semigroups, some results about the rate of the convergence (in the operator norm, too) and further references. A number of references concerning the Chernoff and Trotter-Kato product formulas will be given later.
@@
[Pfei85approximation, Pfei84probabilistic, Pfei86some, Pfei82general] use probabilistic methods to develop a unified exposition of many approximation formulas (see also [BentPau04optimal]).
@@
it holds that (1) $\Rightarrow$ (2) and (2) $\Leftrightarrow$ (3).
@@
A probabilistic formulation of Theorem 4 for Feller processes can be found in [kallenberg]. Trotter-Kato theorems are behind some results about approximations of Feller processes with Markov chains, in particular [kallenberg].
@@
Counterexamples for the Trotter-Kato theorem, in particular of probabilistic origin, can be found in [Bob07limitations]; variational and nonlinear versions of the Trotter-Kato theorem, in [ItoKa02evolution]; generalizations for norm operator topology, in [Zag22notes, CaZag01operator].
The next theorem is called the Chernoff product formula and is another fundamental tool in the approximation theory for $C_{0}-$semigroups.
@@
The formula (3.7) provides an approximation of the semigroup $(T_{t})_{t\geq 0}$ in the terms of the family $V.$ The exposition of the method of obtaining such approximations is given in a survey [But20Method]; further references and applications to stochastic processes including those on manifolds and in domains with Dirichlet and Robin boundary conditions can be found in [But20Method, But18Chernoff, But18Chernoff_for_killed, MaMoReSmo23Chernoff, ButGroSmo10Lagrangian, Nitt09approximations, Sha11Chernoffs, Ne09ChernoffArx, SmoWeizWi07Chernoffs, MaMoReSmo23Chernoff, Ob06representation, SmoToTru02Hamiltonian, But08Feynman, ButSchiSmo11Hamiltonian, ButGroSmo10Lagrangian, But18Chernoff, SmoWeizWi00Brownian]. Other references given later in the text for the Trotter-Kato theorem in the context of the approximation theory are relevant, too.
The following additional references illustrate a possibility to strengthen the conclusion of the Chernoff product formula: [Zag22notes, Zag17comments, Za20notes_chernoff_formula, CaZag01operator, GalRe22rate, GalRe21upper_arx, Prud20speedArx] establish convergence in the operator norm an/or obtain estimates of the rate of the convergence for self-adjoint operators and quasi-sectorial contraction semigroups etc.; [Pau04operator-norm, Zag17comments, Zag22notes, CaZag01operator] obtain estimates by using probabilistic arguments. The convergence can be arbitrary slow and may not hold in a stronger topology.
@@
([But20Method]; cf. [Lejay18Girsanov]) Assume that real-valued functions $a,\sigma$ are Lipschitz continuous and bounded, $\inf_{u\in{\mathbb{R}}}|\sigma(u)|>0$ and consider an SDE
@@
for sufficiently smooth compactly supported $f.$ See also [SmoWeizWi07Chernoffs, MaMoReSmo23Chernoff, SmoWeizWi00Brownian] for other applications of the Chernoff product formula to pathwise approximations of random processes. Note that this result does not recover the first order of the weak Euler-Maruyama scheme.
@@
([ButGroSmo10Lagrangian]) Let $g_{\mu}$ be the one-dimensional Gaussian density with 00 mean and variance $\mu.$ (3.10) can be rewritten as
@@
Such a representation of the original semigroup as the limit of iterated integrals w.r.t. some finite-dimensional projections of a (pseudo)measure on a phase space (the Wiener measure on $C([0;T])$ here) is called the Feynman formula [SmoToTru02Hamiltonian] [^11] The definition of the Feynman formula may slightly vary (c.f. [ButSchiSmo11Hamiltonian]). and presents an alternative approach to the Feynman-Kac formula. See [But20Method, But18Chernoff_for_killed, SmoWeizWi07Chernoffs, MaMoReSmo23Chernoff, Ob06representation, SmoToTru02Hamiltonian, But08Feynman, ButSchiSmo11Hamiltonian, ButGroSmo10Lagrangian, But18Chernoff, SmoWeizWi00Brownian] for other results on Feynman formulas.
@@
and passing to the limit in (3.10) and (12) gives a particular case of the Girsanov theorem (e.g. [LipShi77Statistics]) for $x$ and $y:$
@@
The behavior of a solution of an PDE and its differentiability at $t=0$ (and thus the behavior of the corresponding semigroup) can be a subtle moment in applications, particularly when a probabilistic interpretation is used. A typical example is the function $u(x,t)=\operatorname{E}_{x}f(\xi(\min\{t,\tau\})),$ where $\tau$ is the moment a Feller process $\xi$ hits the boundary of a domain $D.$ If $f=\!\!\,{\textbf{1}}_{\partial D}$ then $u=\operatorname{P}(\tau<t)$ is discontinuous at $t=0$ though $u\in C^{2,1}(D^{int}\times(0;\infty))$ usually. As a result, the weak variational formulation in the terms of Gelfand triples [PaWin08First] or the weak distributional formulation [LeRaRey22probabilistic, TriZab11Pfaffian] can be suggested as an alternative (see also [FeePop15stochastic, Wang17viscosity]).
@@
One particular example of the versatility of the Chernoff product formula is that it can be used to prove the central limit theorem [Goldstein85semigroups] .